The maximum and minimum magnitude of 2i+3j and 5i-2j

Grade:12th pass

1 Answers

Shailendra Kumar Sharma
6 years ago
Two vectors A and B can give the resultant as 
R2=A2+B2+2ABcos(theta)
So the maximum is R=A+B and minimum is A-B
Now 2i+3j and 5i-2j has magnitude of \sqrt{4+9}  =\sqrt{13}}and \sqrt{25+4}  =\sqrt{29}respectively  so maximum will be  A+B=  \sqrt{13}}+\sqrt{29}=8.99 and minimum be 1.78